THE ELECTRONIC SPECTRUM OF H_{2}PO, THE PROTOTYPICAL PHOSPHORYL FREE RADICAL

\underline{MOHAMMED A. GHARAIBEH}, DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

The electronic spectrum of the H_{2}PO radical has been identified by laser-induced fluorescence (LIF) and single vibronic level (SVL) emission techniques. The radical was produced in a pulsed electric discharge jet using a precursor mixture of phosphine (PH_{3}) and carbon dioxide in high-pressure argon and the B̃^{2}A′ - X̃^{2}A′ electronic transition was detected in the 410-338 nm region. Low resolution LIF and SVL emission spectra of H_{2}PO and D_{2}PO have been recorded and the a′ vibrational frequencies have been determined in both states. High-resolution spectra of the 0_{0}^{0} bands of H_{2}PO and D_{2}PO, which consist of strong a-type and weaker c-type components, were recorded. The spectra have been rotationally analyzed and the excited state molecular structure determined. The spectrum of H_{2}PO will be discussed in comparison with the spectra of other phosphoryl and arsenyl radicals that have been recently studied in our laboratory.